OpenStudy (anonymous):

2(x+6)=-2(x-4) can someone solve this? i'm stuck

OpenStudy (anonymous):

x=-1 http://www.algebra.com/services/rendering/simplifier.mpl

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

She solved it using a program, but do you want to learn how?

OpenStudy (anonymous):

2(x+6)=-2(x-4) divide by 2 first. Since both sides have this in common it will be a quick way to simplify the problem. x+6=-(x-4) Now you are going to have to distribute the negative sign, which is really a (-1) to each term inside the ( ). x+6=-x+4 We need to group the common terms, constants with constants, like variables with like variables. To do this we can add something to each side of the equation, thus cancelling it from one side and combining it with the like term on the other. [I'll show this] x+(x)+6-(6)=(x)-x+4-(6) Here I added an x to each side. I chose to do this to keep the x term, what we are solving for, positive. Then I needed to subtract 6 from each side to get rid of it on the left and combine it with the other constant on the right. 2x=-2 We need to get rid of the coefficient in front of x. To do this we can divide both sides by 2. This will get x by itself and leave us with our answer. x = -1